Determine horizontal velocity with which a stone must be projected horizontally from a point P, so that it may hit the inclined plane perpendicularly. the inclination of the plane is 'theta' and point P is at height 'h' above the foot of the incline

Grade:11

1 Answers

Vikas TU
7 years ago
let y be the distance when it hots the inclined plane. 
The height abob=ve plane would be = h – y
then
in horizontal direcn.
vx = v
and in vertical direcn.
vy = root2 * g * y
from the geometry of inclined plane,
tanthetha = vx/vy = v/root(2)*g*y
or
v = root2*g**tan(thetha).